Question 1

Bayou Oil Corporation had the following information and account balances for the years shown relating to Lease No. 1.

12/31/Y1

12/31/Y2

Proved property – cost …………………………………………..

$ 40,000

$ 40,000

Accumulated DD&A – proved property…………………………

4,000

Wells and equipment – IDC ……………………………………..

400,000

600,000

Accumulated DD&A – wells and equipment – IDC …………….

60,000

Wells and equipment – L&WE…………………………………..

300,000

420,000

Accumulated DD&A – wells and equipment – L&WE………….

45,000

Year 1

Year 2

Proved reserves, 12/31 - Oil …………………………..

30,000 bbl

50,000 bbl

                                        Gas………………………….

450,000 Mcf

600,000 Mcf

Proved undeveloped reserves, 12/31 - Oil ……………                                                     

10,000 bbl

12,000 bbl

                                                             Gas……………

200,000 Mcf

120,000 Mcf

Production during 2015 and 2016 - Oil ………………                                           

5,000 bbl

7,000 bbl

                                                         Gas………………

50,000 Mcf

70,000 Mcf

REQUIRED: Compute DD&A for the year ended 12/31/Year 2 using:

  1. A common unit of measure based on equivalent Mcf
  2. Gas as the dominant mineral
  3. Same relative Proportion
Answer #1

Firstly, DD&A is Depreciation, Depletion & Amortization, some of the accounting techniques that companies use to gradually expense resources of economic value.

Computation of Depletion:

Depletion expense = Cost - Salvage Value * Nos of units extracted

estimated nos of units

= $40,000 - 0 * 10,000 = $80,000

5,000

Computation of Amortization:

Amortization = Intangible asset costs / estimated useful life

= 5000/2 = 2500